Calculate the equilibrium constant for each of the reactions at 25∘C∘C.

Standard Electrode Potentials at 25 ∘C
Reduction Half-Reaction
Cl2(g)+2e ---> 2CI- 1.36

I2(s)+2e --> 2I- 0.54

Part A

Cl2(g)+2I−(aq)→2Cl−(aq)+I2(s)

K= ?
Answer #1

from data table:
Eo(I2(s)/2I-) = 0.54 V
Eo(Cl2(g)/2Cl-) = 1.36 V

As per given reaction,
cathode is (Cl2(g)/2Cl-)
anode is (I2(s)/2I-)

Eocell = Eocathode - Eoanode
= (1.36) - (0.54)
= 0.82 V


here, number of electrons being transferred, n = 2
Eo = (2.303*R*T)/(n*F) log Kc
At 25 oC or 298 K, R*T/F = 0.0592
So, Eo = (0.0592/n)*log Kc
0.82 = (0.0592/2)*log Kc
log Kc = 27.7027
Kc = 5.043*10^27
Answer: 5.04*10^27